# Topology-Enabled Switchable Unidirectional Radiative Band in a Bilayer Photonic Crystal

### v1

@@ Abstract
Introduction- A central objective of nanophotonics is to shape how optical modes in open systems couple to the continuum, since this radiative coupling governs energy leakage, quality factors, and the angular distribution of emission [1, 2, 3]. In periodic dielectric structures, this control is encoded in the dispersion and symmetry of guided resonances, which can be engineered through geometry but is ultimately constrained in single-layer platforms by limited degrees of freedom [4, 3, 5, 6, 7, 8, 9]. Multilayer metasurfaces and photonic crystals (PhCs) overcome this restriction by introducing an additional vertical dimension for mode hybridization: stacking spatially separated resonant layers creates a coupled-mode space in which radiative channels can interfere, be redistributed between layers, and acquire new effective symmetries. This enlarged design space has enabled the realization of flatband dispersions[10, 11, 12, 13, 14], enhanced nonlinearities[15, 16], topological effects in synthetic momentum space[17, 18], and polarization singularities[19, 20, 21, 22].
@@ Acknowledgements.
S.T.H, T.T.H.D and V.V acknowledge funding support from A*STAR MTC-Programmatic Fund (M24N9b0122) and National Research Foundation, Singapore – Frontier Competitive Research Grant (NRF-F-CRP-2024-0009). Z. Y., C. D., acknowledge funding support from the National Research Foundation Competitive Research Program (NRF-CRP29-2022-0003), Ministry of Education, Singapore under its AcRF Tier 2 grant (MOE-T2EP50121-0012), and AcRF Tier 1 grant (RG140/23). This project is partly funded by the French National Research Agency (ANR), STRONG-NANO (ANR-22-CE24-0025), POLAROID (ANR-24-CE24-7616-01), 441SUPER-HERO (ANR-25-CE24-4066), and CHOCOLAT (ANR- 44225-CE47-6432)

### v2

@@ Abstract
Introduction- A central objective of nanophotonics is to shape how optical modes in open systems couple to the continuum, since this radiative coupling governs energy leakage, quality factors, and the angular distribution of emission [1, 2, 3]. In periodic dielectric structures, this control is encoded in the dispersion and symmetry of guided resonances, which can be engineered through geometry but is ultimately constrained in single-layer platforms by limited degrees of freedom [4, 3, 5, 6, 7, 8, 9]. Multilayer metasurfaces and photonic crystals (PhCs) overcome this restriction by introducing an additional vertical dimension for mode hybridization: stacking spatially separated resonant layers creates a coupled-mode space in which radiative channels can interfere, be redistributed between layers, and acquire new effective symmetries. This enlarged design space has enabled the realization of flatband dispersions[10, 11, 12, 13, 14], enhanced nonlinearities[15, 16], topological effects in synthetic momentum space[17, 18, 19, 20], and polarization singularities[21, 22, 23, 24].
A particularly appealing consequence of this multilayer interference is the ability to engineer strong directional imbalance between the upward and downward radiation channels. This idea is captured by the recently introduced concept of unidirectional guided resonances (UGRs), which provides a route to mirror-free directional emission in photonic crystal slabs [25, 26, 27]. UGRs arise when the coupling to the two half-spaces becomes strongly imbalanced, producing asymmetric far-field responses that can persist across a photonic band and has now been demonstrated in bilayer metasurfaces [18, 28, 29]. Despite this progress, existing UGR implementations face three recurring limitations: low quality factors due to intrinsically strong radiative leakage, limited post-fabrication reversibility, and directionality predominantly at oblique angles rather than near-normal emission, which constrains their use in high-performance and reconfigurable devices.
@@ Acknowledgements.
S.T.H, T.T.H.D and V.V acknowledge funding support from A*STAR MTC-Programmatic Fund (M24N9b0122) and National Research Foundation, Singapore – Frontier Competitive Research Grant (NRF-F-CRP-2024-0009). Z. Y., C. D., acknowledge funding support from the National Research Foundation Competitive Research Program (NRF-CRP29-2022-0003), Ministry of Education, Singapore under its AcRF Tier 2 grant (MOE-T2EP50121-0012), and AcRF Tier 1 grant (RG140/23). This project is partly funded by the French National Research Agency (ANR), STRONG-NANO (ANR-22-CE24-0025), POLAROID (ANR-24-CE24-7616-01), SUPER-HERO (ANR-25-CE24-4066), and CHOCOLAT (ANR- 44225-CE47-6432)